# On vector-valued functional equations with multiple recursive terms

### v2

@@ 9.2 The joint Laplace-Stieltjes transform
Letting $n\to\infty$, assuming $D_{j}^{(n)}(s_{1},s_{2})\to\tilde{D}_{j}(s_{1},s_{2})$, $K_{i}^{(n)}(s_{1},s_{2})\to\tilde{K}_{i}(s_{1},s_{2})$, $C_{1,1}^{(n)}(s_{1},s_{2})\to\tilde{C}_{1,1}(s_{1},s_{2})$, $C_{1,j}^{(n)}(s_{1},s_{2})\to\tilde{C}_{1,j}(s_{1},s_{2})$, $C_{i,1}^{(n)}(s_{1},s_{2})\to\tilde{C}_{i,1}(s_{1},s_{2})$, and having in mind that $f(0,0)=1$, (97) becomes
@@
$\begin{array}[]{rl}f(s_{1},s_{2})=&\tilde{C}_{1,1}(s_{1},s_{2})A_{1,1}+\sum_{i=2}^{\infty}\tilde{C}_{i,1}(s_{1},s_{2})A_{i,1}+\sum_{j=2}^{\infty}\tilde{C}_{1,j}^{(n)}(s_{1},s_{2})A_{1,j}\vskip 5.69054pt\\ &-\sum_{j=1}^{\infty}\tilde{D}_{j}(s_{1},s_{2})f(0,a_{2}^{j}\lambda_{2})-\sum_{i=1}^{\infty}\tilde{K}_{i}(s_{1},s_{2})f(a_{1}^{i}\lambda_{1},0)+\prod_{k=0}^{\infty}c_{1}(\sigma_{1,1}^{(k)}(s_{1},s_{2})).\end{array}$ (98)
@@
Thus, we have to find $A_{i,j}$, and $f(0,a_{2}^{j}\lambda_{2})$, $f(a_{1}^{i}\lambda_{1},0)$, $i,j=1,2,\ldots$. Note that as the number of iterations increases, i.e., $i\to\infty$, $A_{i,1}:=f(a_{1}^{i}\lambda_{1},a_{2}\lambda_{2})\to f(0,a_{2}\lambda_{2})$, and $j\to\infty$, $A_{1,j}:=f(a_{1}\lambda_{1},a_{2}^{j}\lambda_{2})\to f(a_{1}\lambda_{1},0)$ thus in practice, we need to obtain a few terms $A_{i,1}$, $A_{1,j}$, $i,j\geq 1$, and after that, all the terms converge to $f(0,a_{2}\lambda_{2})$, $f(a_{1}\lambda_{1},0)$.
@@ 10 Conclusion and suggestions for future research
In this work, we dealt with vector-valued recursions between random vectors that lead to functional equations of the form (1), (2). In Section 2, we cope with the transient analysis of a Markov-modulated $M/G/1$-type reflected autoregressive process in which a vector-valued functional equation of the type (1) naturally arises. In Sections 3-5 we dealt with the stationary behavior of vector-valued recursions that arise from queueing processes and autoregressive processes where dependencies among random variables are dictated by (32), (39), and for which a similar vector-valued functional equation arises; see (2). In Section 6, we also cope with a queueing model, where dependencies are based on (32), but the vector-valued functional equation is somehow different compared to the others treated in this work, in the sense that the involved mappings $a_{i}(s)=s+\mu_{i}$, $i=1,\ldots,N$ are commutative, but there are not contractions. However, even in that case we were able to solve it iteratively. Section 7 is devoted to the analysis of a modulated 2-queue ASIP network with consumption, which led to an even more general functional equation. In Section 8, we focused on the modulated integer-valued reflected autoregressive process. Finally, in Section 9, we considered a reflected VAR(1) process, where the autoregressive parameter is now a given $N\times N$ matrix.
An interesting topic for future research is to cope with the case where we have vector-valued functional equations in which the involved contraction mappings are not commutative. For the scalar case (i.e., non-modulated), there exist some available scarce results, see [7], [29]. It would be interesting to further investigate what can still be accomplished both for the scalar and the vector-valued case, when we are dealing with a noncommutative contraction mapping. Other options for future research refer to the case where $R_{n}(X_{n})$ may take negative values.

### v3

@@ 4.2.2 The joint Laplace-Stieltjes transform
Note that each term appearing in the recursive definitions of $C_{i,j}^{(n)}(s_{1},s_{2})$, $D_{j}^{(n)}(s_{1},s_{2})$, and $K_{i}^{(n)}(s_{1},s_{2})$ is a finite product involving factors of the form $c_{1}(\cdot)$ and $c_{\ell}(\cdot)$, $\ell=2,3,4$, evaluated at iterated arguments of the form $(a_{1}^{k}s_{1},a_{2}^{k}s_{2})$ or mixed arguments where one component is fixed (e.g., $(a_{1}^{k}s_{1},a_{2}^{j}\lambda_{2})$).
The main result is summarized in the following theorem.
Theorem 32
Consider the functional equation (105) and assume that the functions $c_{k}(s_{1},s_{2})$, $k=1,2,3,4$ defined in (104) with $c_{1}(0,0)=1$, $c_{2}(0,0)=c_{3}(0,0)=c_{4}(0,0)=0$. Let $a_{1},a_{2}\in(0,1)$ and define $a^{*}:=\max\{a_{1},a_{2}\}<1$. Then, for all $(s_{1},s_{2})$ with $Re(s_{1})\geq 0$, $Re(s_{2})\geq 0$, the solution of (105) is given by
@@
$\begin{array}[]{rl}f(s_{1},s_{2})=&\tilde{C}_{1,1}(s_{1},s_{2})A_{1,1}+\sum_{i=2}^{\infty}\tilde{C}_{i,1}(s_{1},s_{2})A_{i,1}+\sum_{j=2}^{\infty}\tilde{C}_{1,j}(s_{1},s_{2})A_{1,j}\vskip 5.69054pt\\ &-\sum_{j=1}^{\infty}\tilde{D}_{j}(s_{1},s_{2})f(0,a_{2}^{j}\lambda_{2})-\sum_{i=1}^{\infty}\tilde{K}_{i}(s_{1},s_{2})f(a_{1}^{i}\lambda_{1},0)+\prod_{k=0}^{\infty}c_{1}(\sigma_{1,1}^{(k)}(s_{1},s_{2})),\end{array}$ (107)
where the coefficients $\tilde{C}_{i,j}(s_{1},s_{2})$, $\tilde{D}_{j}(s_{1},s_{2})$, and $\tilde{K}_{i}(s_{1},s_{2})$ are the limits of the recursively defined sequences $C_{i,j}^{(n)}(s_{1},s_{2})$, $D_{j}^{(n)}(s_{1},s_{2})$, and $K_{i}^{(n)}(s_{1},s_{2})$, respectively. Moreover, all series and the infinite product in (107) converge absolutely.
Proof. Iterate (105) n1n-1 times to obtain (106), where all coefficients are defined recursively. Since $a_{1},a_{2}\in(0,1)$, we have $\sigma_{1,1}^{(k)}(s_{1},s_{2})\to(0,0)$ as $k\to\infty$. Since $c_{1}(s_{1},s_{2})$ is analytic close to $(0,0)$ and $c_{1}(0,0)=1$, there exists $C>0$ such that $|c_{1}(s_{1},s_{2})-1|\leq C(|s_{1}|+|s_{2}|)$. Thus, for $(s_{1},s_{2})$ sufficiently close to $(0,0)$ we have
$|c_{1}(\sigma_{1,1}^{(k)}(s_{1},s_{2}))-1|\leq C(a^{*})^{k}(|s_{1}|+|s_{2}|).$
Since $\sum_{k=0}^{\infty}(a^{*})^{k}<\infty$, it follows that
$\sum_{k=0}^{\infty}|c_{1}(\sigma_{1,1}^{(k)}(s_{1},s_{2}))-1|<\infty,$
and therefore $\prod_{k=0}^{\infty}c_{1}(\sigma_{1,1}^{(k)}(s_{1},s_{2}))$, converges absolutely. Having in mind that $f(0,0)=1$, we have $f(a_{1}^{n}s_{1},a_{2}^{n}s_{2})\to 1$. Thus, as $n\to\infty$
$\prod_{k=0}^{n-1}c_{1}(\sigma_{1,1}^{(k)}(s_{1},s_{2}))f(a_{1}^{n}s_{1},a_{2}^{n}s_{2})\to\prod_{k=0}^{\infty}c_{1}(\sigma_{1,1}^{(k)}(s_{1},s_{2})).$
All recursively defined quantities $C_{i,j}^{(n)}(s_{1},s_{2})$, $D_{j}^{(n)}(s_{1},s_{2})$, and $K_{i}^{(n)}(s_{1},s_{2})$ are finite sums of products involving factors $c_{1}(\cdot)$ and $c_{l}(\cdot)$, $l=2,3,4$, evaluated at iterated arguments of the form $(a_{1}^{k}s_{1},a_{2}^{k}s_{2})$ or at mixed arguments where one component is fixed (e.g., $(a_{1}^{k}s_{1},a_{2}^{j}\lambda_{2})$). In particular, products of the form
$\prod_{m=0}^{k-1}c_{1}(\sigma_{1,1}^{(m)}(s_{1},s_{2})),$
converge to a finite limit, and therefore are uniformly bounded. The same holds for products of $c_{1}(\cdot)$ evaluated at the corresponding iterated arguments appearing in the recursion. Since $c_{2}(0,0)=c_{3}(0,0)=c_{4}(0,0)=0$ and the functions are continuous, there exists $C>0$ such that for $(s_{1},s_{2})$ sufficiently close to $(0,0)$,
$|c_{l}(s_{1},s_{2})|\leq C(|s_{1}|+|s_{2}|),\,l=2,3,4.$
Hence,
$|c_{l}(\sigma_{1,1}^{(k)}(s_{1},s_{2}))|\leq C(a^{*})^{k}(|s_{1}|+|s_{2}|),\,l=2,3,4.$
Hence, each term appearing in the recursive definitions is bounded by a geometric sequence of order $(a^{*})^{k}$, and therefore all corresponding series converge absolutely. The fact that $c_{2}(0,0)=c_{3}(0,0)=c_{4}(0,0)=0$ ensures that the corresponding terms vanish at the fixed point $(0,0)$. This implies that these terms introduce additional decay along the iterations, which guarantees the absolute convergence of the series in (107). In this sense, $c_{1}(s_{1},s_{2})$ governs the main recursive structure, while $c_{2}(s_{1},s_{2})$, $c_{3}(s_{1},s_{2})$, $c_{4}(s_{1},s_{2})$ act as vanishing perturbations (this is analogous to the case $\tilde{V}(0)=\tilde{0}$ in Theorem 1).
Moreover, the recursive increments can be written as finite sums of terms involving such products evaluated at increasing indices. Each increment can be written as a finite sum of terms of the form
$\left(\prod_{m=0}^{k-1}c_{1}(\sigma_{1,1}^{(m)}(s_{1},s_{2}))\right)c_{l}(\sigma_{1,1}^{(k)}(s_{1},s_{2})),\,l=2,3,4,$
which are bounded by $C(a^{*})^{k}$. Since $\sum_{k=0}^{\infty}(a^{*})^{k}<\infty$, the increments are summable and therefore tend to zero. Therefore, the sequences $C_{i,j}^{(n)}(s_{1},s_{2})$, $D_{j}^{(n)}(s_{1},s_{2})$, and $K_{i}^{(n)}(s_{1},s_{2})$ are Cauchy and converge. Finally, since (106) holds for all $n$ and all terms on the right-hand side converge as $n\to\infty$, by taking the limit as $n\to\infty$ in (106) we obtain (107).
@@
We finally have to find $A_{i,j}$, and $f(0,a_{2}^{j}\lambda_{2})$, $f(a_{1}^{i}\lambda_{1},0)$, $i,j=1,2,\ldots$. Note that as the number of iterations increases, i.e., $i\to\infty$, $A_{i,1}:=f(a_{1}^{i}\lambda_{1},a_{2}\lambda_{2})\to f(0,a_{2}\lambda_{2})$, and $j\to\infty$, $A_{1,j}:=f(a_{1}\lambda_{1},a_{2}^{j}\lambda_{2})\to f(a_{1}\lambda_{1},0)$ thus in practice, we need to obtain a few terms $A_{i,1}$, $A_{1,j}$, $i,j\geq 1$, and after that, all the terms converge to $f(0,a_{2}\lambda_{2})$, $f(a_{1}\lambda_{1},0)$.
@@ 5 Conclusion and suggestions for future research
In this work, we dealt with vector-valued recursions between random vectors that lead to functional equations of the form (1), (2), (4), (5). The general theory to treat (1), (2) is presented in Section 2. The theoretical results are applied in a series of queueing related examples in Section 3. In Section 4, we dealt with some multidimensional processes that result in (4), (5). Although, their form is different from that of (1), the solution machinery is similar. In particular, we focused on a modulated ASIP two queue tandem system and on a reflected VAR(1) process, where the autoregressive parameter is now a given $N\times N$ matrix.
An interesting topic for future research is to cope with the case where we have vector-valued functional equations in which the involved contraction mappings are not commutative. For the scalar case (i.e., non-modulated), there exist some available scarce results, see [7], [29]. It would be interesting to further investigate what can still be accomplished both for the scalar and the vector-valued case, when we are dealing with a noncommutative contraction mapping. Other options for future research refer to the case where $R_{n}(Y_{n})$ may take negative values, as well as to fully consider the case where the autoregressive parameter is a scalar or a random matrix. A first attempt was made in subsection 4.2, where the autoregressive parameter was a scalar matrix, and where our focus was on the case where this matrix is diagonal, and specifically when $N=2$. A natural extension should be to investigate the case where this matrix is of arbitrary dimension, and not only diagonal.
@@ Acknowledgements
Appendix A Proof of Theorem 3
By iterating (4) $n$ times, we obtain,
$\tilde{Z}(s,t)=\sum_{k=0}^{n-1}\sum_{\gamma\in\Gamma_{k}}\mathcal{K}_{\gamma}(s,t)\,K(s_{k},t_{k})+\sum_{\gamma\in\Gamma_{n}}\mathcal{K}_{\gamma}(s,t)\,\tilde{Z}(s_{n},t_{n}).$ (114)
Due to the commutative contraction mappings $\alpha_{i}(u)$ with common fixed point $a$, we know that there exists $\kappa\in(0,1)$ such that for $u\in\mathbb{C}_{+}$, $|\alpha_{i}(u)-a|\leq\kappa|u-a|$. By the contraction property and the fact that $\alpha_{i}(a)=a$, we obtain recursively
$|t_{k}-a|=|\alpha_{i_{k}}(t_{k-1})-\alpha_{i_{k}}(a)|\leq\kappa|t_{k-1}-a|\leq\kappa^{k}|t_{0}-a|=\kappa^{k}|t-a|.$
Since $s_{m}$ is either $s_{m-1}$ or $t_{m-1}$, it follows that $(s_{k},t_{k})\to(a,a)$ as $k\to\infty$.
Step 1: Lipschitz continuity of $\tilde{Z}(s,t)$.
Let $(s,t),(u,v)\in\mathbb{C}_{+}\times\mathbb{C}_{+}$. Then,
$|z_{i}(s,t)-z_{i}(u,v)|$ $=\left|\mathbb{E}\Big(\big(e^{-sW_{1}-tW_{2}}-e^{-uW_{1}-vW_{2}}\big)1_{\{Y=i\}}\Big)\right|\leq\mathbb{E}\Big(\big|e^{-sW_{1}-tW_{2}}-e^{-uW_{1}-vW_{2}}\big|\,1_{\{Y=i\}}\Big).$
Note that
$e^{-sW_{1}-tW_{2}}-e^{-uW_{1}-vW_{2}}$ $=\big(e^{-sW_{1}}-e^{-uW_{1}}\big)e^{-tW_{2}}+e^{-uW_{1}}\big(e^{-tW_{2}}-e^{-vW_{2}}\big).$
Hence,
$|e^{-sW_{1}-tW_{2}}-e^{-uW_{1}-vW_{2}}|$ $\leq|e^{-sW_{1}}-e^{-uW_{1}}|\,|e^{-tW_{2}}|+|e^{-uW_{1}}|\,|e^{-tW_{2}}-e^{-vW_{2}}|.$
Since $Re(s)\geq 0$, $Re(u)\geq 0$, $Re(t)\geq 0$, $Re(v)\geq 0$, it is readily seen that $|e^{-sW_{1}}|\leq 1$, $|e^{-uW_{1}}|\leq 1$, $|e^{-tW_{2}}|\leq 1$, $|e^{-vW_{2}}|\leq 1$. Moreover, by the mean value theorem, we obtain
$|e^{-sW_{1}}-e^{-uW_{1}}|\leq|s-u|\,W_{1},\hskip 18.49988pt|e^{-tW_{2}}-e^{-vW_{2}}|\leq|t-v|\,W_{2}.$
Combining the above we have
$|z_{i}(s,t)-z_{i}(u,v)|$ $\leq\mathbb{E}\big(|s-u|W_{1}\,1_{\{Y=i\}}\big)+\mathbb{E}\big(|t-v|W_{2}\,1_{\{Y=i\}}\big)$
$=|s-u|\,\mathbb{E}(W_{1}1_{\{Y=i\}})+|t-v|\,\mathbb{E}(W_{2}1_{\{Y=i\}}).$
Define
$C_{1}:=\max_{1\leq i\leq N}\mathbb{E}(W_{1}1_{\{Y=i\}}),\hskip 18.49988ptC_{2}:=\max_{1\leq i\leq N}\mathbb{E}(W_{2}1_{\{Y=i\}}).$
Then,
$|z_{i}(s,t)-z_{i}(u,v)|\leq C_{1}|s-u|+C_{2}|t-v|.$
Taking the maximum over $i=1,\ldots,N$, we obtain
$\|\tilde{Z}(s,t)-\tilde{Z}(u,v)\|_{\infty}=\max_{1\leq i\leq N}|z_{i}(s,t)-z_{i}(u,v)|\leq C_{1}|s-u|+C_{2}|t-v|.$
Setting $C:=\max\{C_{1},C_{2}\}$, we conclude that
$\|\tilde{Z}(s,t)-\tilde{Z}(u,v)\|_{\infty}\leq C\big(|s-u|+|t-v|\big),$
which shows that $\tilde{Z}(s,t)$ is Lipschitz continuous on $\mathbb{C}_{+}\times\mathbb{C}_{+}$.
Step 2: Bound on $\mathcal{K}_{\gamma}(s,t)$.
Since $\tilde{P}^{(i)}=e_{i}e_{i}^{T}$, right multiplication by $\tilde{P}^{(i)}$ retains only the $i$-th column of a matrix. Therefore, by induction on $k$, each product $\mathcal{K}_{\gamma}(s,t)$ has at most one nonzero column (namely the $i_{k}$-th column). Let $M(a):=\max\{\|R(a,a)\|_{1},\|T(a)\|_{1}\}$. By analyticity of $R(s,t)$ and $T(s)$, for every $\epsilon>0$ there exists $\delta>0$ such that for $(u,v)$ sufficiently close to $(a,a)$, we have $\|R(u,v)\|_{1},\ \|T(u)\|_{1}\leq M(a)+\epsilon$.
Since $(s_{m},t_{m})\to(a,a)$, for sufficiently large $m$ there exists $k_{0}$ such that for all $m\geq k_{0}$, $(s_{m},t_{m})$ lies in a neighborhood of $(a,a)$. Hence,
$\|\mathcal{K}_{\gamma}(s,t)\|_{1}\leq C_{0}(M(a)+\epsilon)^{k-k_{0}}\leq C(M(a)+\epsilon)^{k},$
for some constant $C>0$ independent of $\gamma$.
Step 3: Decomposition of the remainder.
Let
$R_{n}(s,t):=\sum_{\gamma\in\Gamma_{n}}\mathcal{K}_{\gamma}(s,t)\,\tilde{Z}(s_{n},t_{n}),$
and write $\tilde{Z}(s_{n},t_{n})=\tilde{Z}(a,a)+\big(\tilde{Z}(s_{n},t_{n})-\tilde{Z}(a,a)\big)$, so that
$R_{n}(s,t)=S_{n}(s,t)+D_{n}(s,t),$
where
$\begin{array}[]{rl}S_{n}(s,t):=&\sum_{\gamma\in\Gamma_{n}}\mathcal{K}_{\gamma}(s,t)\,\tilde{Z}(a,a),\\ D_{n}(s,t):=&\sum_{\gamma\in\Gamma_{n}}\mathcal{K}_{\gamma}(s,t)\big(\tilde{Z}(s_{n},t_{n})-\tilde{Z}(a,a)\big).\end{array}$
Step 4: Vanishing of $D_{n}(s,t)$.
By exploiting the projection property and the bounds above, we obtain
$\|D_{n}(s,t)\|_{\infty}\leq\max_{\gamma\in\Gamma_{n}}\|\mathcal{K}_{\gamma}(s,t)\|_{1}\|\tilde{Z}(s_{n},t_{n})-\tilde{Z}(a,a)\|_{\infty}.$
Hence,
$\|D_{n}(s,t)\|_{\infty}\leq C(M(a)+\epsilon)^{n}\kappa^{n}.$
Since $M(a)\kappa<1$, there exists $\epsilon>0$ such that $(M(a)+\epsilon)\kappa<1$, and thus $D_{n}(s,t)\to 0$.
Step 5: Convergence of $(S_{n})$.
Using the recursive structure of $\mathcal{K}_{\gamma}(s,t)$, we obtain (see also the corresponding step in the proof of Theorem 1)
$S_{n+1}-S_{n}=\sum_{\gamma\in\Gamma_{n}}\mathcal{K}_{\gamma}(s,t)\,\big(A_{\gamma}(s_{n},t_{n})-A_{\gamma}(a,a)\big)\tilde{Z}(a,a),$
where
[…]
Since $R(s,t)$ and $T(s)$ are analytic, they are continuously differentiable in a neighborhood of $(a,a)$. Therefore, there exists $L>0$ such that for $(u,v)$ sufficiently close to $(a,a)$,
$\|A_{\gamma}(u,v)-A_{\gamma}(a,a)\|_{1}\leq L(|u-a|+|v-a|),$
uniformly for $A_{\gamma}\in\{R,T\}$. Applying this bound with $(u,v)=(s_{n},t_{n})$ and using the contraction property, we obtain
$\|A_{\gamma}(s_{n},t_{n})-A_{\gamma}(a,a)\|_{1}\leq C\kappa^{n},$
for some constant $C>0$. Then, by using the one-column structure of $\mathcal{K}_{\gamma}(s,t)$ that is inherited by the projection property, we have
$\|S_{n+1}(s,t)-S_{n}(s,t)\|_{\infty}\leq\max_{\gamma\in\Gamma_{n}}\|\mathcal{K}_{\gamma}(s,t)\|_{1}\cdot\|A_{\gamma}(s_{n},t_{n})-A_{\gamma}(a,a)\|_{1}\cdot\|\tilde{Z}(a,a)\|_{\infty}.$
Thus,
$\|S_{n+1}(s,t)-S_{n}(s,t)\|_{\infty}\leq C(M(a)+\epsilon)^{n}\kappa^{n}.$
Since $(M(a)+\epsilon)\kappa<1$, it follows that
$\sum_{n=0}^{\infty}\|S_{n+1}-S_{n}\|_{\infty}<\infty,$
and therefore $(S_{n})$ is a Cauchy sequence in $\mathbb{C}^{N}$, and therefore, it is convergent.
Step 6: Absolute convergence of the series.
Since $(s_{k},t_{k})\to(a,a)$ and $K$ is analytic, $\|K(s_{k},t_{k})\|_{\infty}$ is bounded. Using again the one-column structure due to the projection property, we obtain
$\left\|\sum_{\gamma\in\Gamma_{k}}\mathcal{K}_{\gamma}(s,t)\,K(s_{k},t_{k})\right\|_{\infty}\leq C(M(a)+\epsilon)^{k}.$
Since $M(a)\kappa<1$, there exists $\epsilon>0$ such that $(M(a)+\epsilon)\kappa<1$, and the series converges absolutely.
Letting $n\to\infty$ in (114), we finally obtain
$\tilde{Z}(s,t)=\sum_{k=0}^{\infty}\sum_{\gamma\in\Gamma_{k}}\mathcal{K}_{\gamma}(s,t)\,K(s_{k},t_{k})+\lim_{n\to\infty}\sum_{\gamma\in\Gamma_{n}}\mathcal{K}_{\gamma}(s,t)\,\tilde{Z}(s_{n},t_{n}),$
which completes the proof.